# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

### v1

@@ 3.1 Algorithm
       $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\big{(}\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T\big{)},\pi^{\prime}_{t}=\pi_{t-1}$;
        // Update policy with RL
4       $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$ with $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}_{t}}$;
       $y_{t}=\mathbbm{1}(\tau_{t}\succ\tau^{\prime}_{t})$;
@@
5       $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\left({y_{t},x_{t}}\right)$;
6       $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$;
       $\hat{\theta}_{t+1}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t+1}}(\theta)$ ;
@@ 6       $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$;
Output: Policy $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\bar{\theta}_{T},\varepsilon,\delta^{\prime}}\right)$ with $\bar{\theta}_{T}=\frac{1}{T}\sum_{t=1}^{T}\hat{\theta}_{t}$.
Algorithm 1 Randomized Preference Optimization (RPO)
@@ 3.2 Theoretical results
We analyze the regret of Algorithm 1 and the suboptimality of its output policy $\hat{\pi}$.
@@ 3.2.1 Regret analysis
We show that Algorithm 1 with $\alpha=1$ has sublinear regret with high probability.
Theorem 3.4.
Using Algorithm 1 with $\alpha=1$, it holds for any $\delta>0$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
$R(T)\leq\mathcal{O}\left(\sqrt{\kappa d^{3}T\log(dT/\delta)^{3}}\right).$
The regret bound of Theorem 3.4 matches the best existing bounds for algorithms with randomized exploration in reinforcement learning, see (Efroni et al., 2021; Ouhamma et al., 2023). In addition, due to learning from preferences, we have an extra $\sqrt{\kappa}$ factor (see Remark 3.2), which is in line with other recent work on RLHF (Wu and Sun, 2023; Das et al., 2024).
@@
For episodic tabular MDPs Pacchiano et al. (2021) prove a regret bound of $\widetilde{\mathcal{O}}(\kappa d\sqrt{T})$. Similarly, Chen et al. (2022) considers episodic linear MDPs and derives a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{HT})$, avoiding dependence on $\kappa$ by assuming a linear preference model. However, these approaches rely on a type of optimism which is computationally intractable (see Appendix E.2). Similar to us, Wu and Sun (2023) avoids this challenge by resorting to randomized exploration and proves a $\widetilde{\mathcal{O}}(d^{3}\sqrt{\kappa T})$ regret bound for linear MDPs. Compared to Wu and Sun (2023), our analysis improves the dependence on dimensionality from $d^{3}$ to $d^{3/2}$ and avoids the need for truncation techniques on the value function. Furthermore, our settings differ in two key points: First, we assume access to an RL oracle without restricting the class of MDPs, whereas Wu and Sun (2023) considers linear MDPs. Second, their approach is model-based, while Algorithm 1 is oracle-based and can accommodate both model-based and model-free implementations.
@@
The full proof of this Theorem is provided in Appendix B. Analogous to the analysis of linear Thompson sampling (Abeille and Lazaric, 2017), the main idea is to control the regret by showing that randomized exploration ensures a constant probability of optimism. However, compared to the linear bandit analysis, our setting comes with additional challenges: First, due to preference-based learning we require a different regret decomposition accounting for the reference policy. Second, as we observe preference feedback on trajectories rather than policies, we need to apply Freedman’s inequality (see Lemma A.9) to control the deviation between expected and observed features. Lastly, as we assume a PAC RL oracle – in place of an exact maximization oracle – we need to carefully track the resulting approximation error.
@@ 3.2.2 Suboptimality gap
The pure exploration version of RPO with $\alpha=0$, in contrast to $\alpha=1$, may incur linear regret during learning as the policy $\pi_{t}$ can be highly suboptimal. However, Theorem 3.5 below shows that it outputs a policy that is $\tilde{\mathcal{O}}(1/\sqrt{T})$-optimal.
Theorem 3.5.
Using Algorithm 1 with $\alpha=0$, it holds for any $\delta>0$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
@@
Except for the extra $\sqrt{d}$ dependency, which is inherent to approaches based on Thompson sampling, [^3] Recently, Abeille et al. (2025) show that for certain classes of linear bandits the additional $\sqrt{d}$ factor can be avoided. However, their assumptions are not directly applicable to our preference-based RL setting. we match the last iterate guarantee proposed by Das et al. (2024) for contextual linear bandits, but with an algorithm that is tractable in the full RL setting.
@@
Few works provide suboptimality guarantees in preference-based RL. Wang et al. (2023) propose a meta-algorithm interfacing with a PAC-RL oracle that outputs an $\varepsilon$-optimal policy after $\widetilde{\mathcal{O}}(\kappa^{2}d^{3}/\varepsilon^{2})$ queries. A different approach by Zhan et al. (2024) leverages optimal design to prove a bound of $\widetilde{\mathcal{O}}((|S|^{2}|A|d+\kappa^{2}d^{2})/\varepsilon^{2})$, but their method relies on an intractable maximization oracle. In comparison, our algorithm achieves a bound of $\widetilde{\mathcal{O}}(\kappa d^{3}/\varepsilon^{2})$, improving the dependence on $\kappa$ over both prior results. The additional factor of $d$ compared to Zhan et al. (2024) is expected, as our method is randomized rather than optimistic, see (Abeille and Lazaric, 2017). Notably, to the best of our knowledge, this is the first algorithm, whether in preference-based or reward-based RL, that provides last-iterate guarantees using randomized exploration.
@@
The proof of Theorem 3.5, presented in Appendix C, builds on Das et al. (2024)’s suboptimality analysis for the contextual bandits setting. However, to sidestep the intractability of maximizing an exploration bonus over policies, we leverage randomized exploration (Abeille and Lazaric, 2017) to ensure a constant probability of optimism. This allows us to derive a bound on the output policy’s suboptimality that mirrors the regret bound, without needing additional assumptions.
@@ 3.3 Practical limitations
While Algorithm 1 is tractable and statistically efficient, it presents certain limitations. First, invoking an RL oracle at every round (line 5) can cause high latency, especially in continuous state spaces. Second, issuing preference queries at each round (line 7) is impractical when the oracle is a human, due to the need for continuous feedback. Finally, requesting a label for all trajectory pairs can be expensive and inefficient, as many comparisons are uninformative.
The next section presents a refined algorithm addressing these limitations. The new approach decouples trajectory collection from query selection, and queries only the most informative comparisons.
@@ 4.1 Algorithm
Input: Number iterations $T$, confidence $\delta>0$, regularization $\lambda>1$, constant $C>0$.
1 Set $\varepsilon=1/\sqrt{T}$, $\delta^{\prime}=\delta/5$, $t_{\operatorname{stop}}=1$.
2 Initialize design matrices $V_{0}=W_{0}=\lambda I$, preference datasets $\mathcal{D}=\mathcal{D}_{0}=\emptyset$, parameter $\hat{\theta}_{0}$.
3 for $t=1,2,\dots,T$ do
4       if $\det(W_{t})>(1+C)\det(W_{t_{\operatorname{stop}}})$ then
             $\mathcal{D}_{\text{opt}},V_{t}=\operatorname{D-OptDes}(\mathcal{D},V_{t_{\operatorname{stop}}},\det(W_{t}))$ ;
              // Greedy D-Optimal design
5             $\mathcal{D}_{t}=\mathcal{D}_{t_{\operatorname{stop}}}$;
@@
                   $x=\phi(\tau)-\phi(\tau^{\prime}),y=\mathbbm{1}(\tau\succ\tau^{\prime})$ ;
                    // Preference feedback
7                   $\mathcal{D}_{t}=\mathcal{D}_{t}\cup\{(x,y)\}$;
@@
            $\hat{\theta}_{t}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t}}(\theta)$ ;
@@
9             $t_{\operatorname{stop}}=t$, $\mathcal{D}=\emptyset$, $\pi^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\hat{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$;
@@
      $\tilde{\theta}_{t}\sim\mathcal{N}(\hat{\theta}_{t_{\operatorname{stop}}},\beta_{t_{\operatorname{stop}}}(\delta^{\prime})^{2}V_{t_{\operatorname{stop}}}^{-1})$;
@@
       $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$;
@@
12       $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$ with $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}}$;
13       If $t=t_{\operatorname{stop}}$, then $W_{t+1}=V_{t}+x_{t}x_{t}^{\top}$, else: $W_{t+1}=W_{t}+x_{t}x_{t}^{\top}$;
14       $\mathcal{D}=\mathcal{D}\cup\{(\tau_{t},\tau^{\prime}_{t})\}$;
15
@@ 14       $\mathcal{D}=\mathcal{D}\cup\{(\tau_{t},\tau^{\prime}_{t})\}$;
Algorithm 2 Lazy Randomized Preference Optimization with Optimal Design (LRPO-OD)
@@
We use an idea from Abbasi-Yadkori et al. (2011) to collect many trajectory pairs without querying the preference oracle. The modification compared to Algorithm 1 is collecting trajectories without updating the MLE $\hat{\theta}_{t_{\operatorname{stop}}}$ until the information gain, represented by $\det(V_{t})$, increases by a multiplicative constant; see line 4 of Algorithm 2. We show that this procedure limits the number of batches to $\mathcal{O}(\log(T))$. In other words, the average (over batches) size of a given batch is of order $\mathcal{O}(T/\log(T))$. A key advantage of this lazy update structure is that the preference queries (line 8) can be collected in parallel across all trajectory pairs within a batch. This significantly reduces the workload of the preference oracle, e.g. a human annotator, by eliminating the need for round-by-round feedback. In particular, the algorithm no longer pauses at each timestep to wait for preference labels. To our knowledge, this is the first approach to support parallelization of queries in online preference-based reinforcement learning, even in settings as simple as dueling bandits.
@@
Due to the submodularity of the $\log\det$ function for $\lambda$ greater than one, the greedy procedure of Algorithm 3 achieves an $(1-1/e)$-approximation to the optimal solution; see (Nemhauser et al., 1978; Krause et al., 2008) and Appendix E.1.
Another key feature of Algorithm 3 is that the while loop is stopped early if $\det(V)$ exceeds the threshold value. This threshold is set as the determinant of the naive design, where every trajectory pair is queried once. When this early termination is satisfied, Algorithm 2 requires fewer samples than Algorithm 1. In particular, since the optimal design maximizes the information gain (measured by $\det(V)$), this termination condition is expected to be satisfied frequently in practice.
Input: Dataset $\mathcal{D}$, current design matrix $V$, threshold for the determinant value $\alpha$.
1 Initialize and dataset $\mathcal{D}_{\text{opt}}=\emptyset$
2 while $\det(V)<\alpha$ and $|\mathcal{D}_{\text{opt}}|\leq|\mathcal{D}|$ do

### v2

@@ 3.1 Algorithm
 $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\big(\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T\big),\pi^{\prime}_{t}=\pi_{t-1}$
 // RL update
3 $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}_{t}}$
4 $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$, $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$
 $y_{t}=\mathbbm{1}(\tau_{t}\succ\tau^{\prime}_{t})$, $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\left\{{\left({x_{t},y_{t}}\right)}\right\}$
@@ 4 $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$, $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$
 $\hat{\theta}_{t+1}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t+1}}(\theta)$
@@
Algorithm 1 RPO–Regret (online preference learning for regret minimization)
Input: $T\in\mathbb{N},\ \delta\in(0,1),\ \lambda>0$
1 Initialize: $\varepsilon=1/\sqrt{T}$; $\delta^{\prime}=\delta/5$; $V_{1}=\lambda I$; $\mathcal{B}=\mathcal{D}=\emptyset$; and choose $\pi_{0}$.
2for $t=1,2,\dots,T$ do
 $\tilde{\theta}_{t}\sim\mathcal{N}\big(0,V_{t}^{-1}\big)$
 // Reward sampling
 $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\big(\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T\big),\ \pi^{\prime}_{t}=\pi_{t-1}$
 // RL update
3 $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\ \tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}_{t}}$
4 $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$, $V_{t+1}=V_{t}+x_{t}x_{t}^{\top}$
 $\mathcal{B}=\mathcal{B}\cup\{(\tau_{t},\tau^{\prime}_{t},x_{t})\}$
 // Defer preference feedback
5
6 end for
7foreach $(\tau,\tau^{\prime},x)\in\mathcal{B}$ do
 $y=\mathbbm{1}(\tau\succ\tau^{\prime})$; $\mathcal{D}=\mathcal{D}\cup\{(x,y)\}$
 // Preference feedback
8
9 end foreach
Output: Policy $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\hat{\theta},\varepsilon,\delta^{\prime}}\right)$ with $\hat{\theta}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}}(\theta)$
// One MLE at end
10
Algorithm 2 RPO–Explore (preference-free exploration and batched reward estimation)
@@ 3.2 Theoretical results
We analyze the regret of RPO-Regret and the suboptimality of the output policy $\hat{\pi}$ of RPO-Explore.
@@ 3.2.1 Regret analysis
We show that RPO-Regret incurs sublinear regret with high probability. {restatable}theoremregret For any $\delta\in(0,1)$ and $T\in\mathbb{N}$, Algorithm 1 satisfies, with probability at least $1-\delta$,
$R(T)=\mathcal{O}\left(\sqrt{\kappa d^{3}T\log(dT/\delta)^{3}}\right).$
The regret bound of Theorem 3.2.1 matches the best existing bounds for algorithms with randomized exploration in reinforcement learning, see (efroni2021reinforcement; ouhamma2023bilinear). In addition, due to learning from preferences, we have an extra $\sqrt{\kappa}$ factor (see Remark 3.1), which is in line with other recent work on RLHF (wu2023making; das2024active).
@@
For episodic tabular MDPs pacchiano2021dueling prove a regret bound of $\widetilde{\mathcal{O}}(\kappa d\sqrt{T})$. Similarly, chen2022human considers episodic linear MDPs and derives a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{HT})$, avoiding dependence on $\kappa$ by assuming a linear preference model. However, these approaches rely on a type of optimism which is computationally intractable in this setting (see Appendix E.2). Similar to us, wu2023making avoid this challenge by resorting to randomized exploration and proves a $\widetilde{\mathcal{O}}(d^{3}\sqrt{\kappa T})$ regret bound for linear MDPs. Compared to wu2023making, our analysis improves the dependence on dimensionality from $d^{3}$ to $d^{3/2}$ and avoids the need for truncation techniques on the value function. Furthermore, our settings differ in two key points: First, we assume access to an RL oracle without restricting the class of MDPs, whereas wu2023making considers linear MDPs. Second, their approach is model-based, while Algorithm 1 is oracle-based and can accommodate both model-based and model-free implementations.
@@
The full proof of Theorem 3.2.1 is provided in Appendix A. Analogous to the analysis of linear Thompson sampling (pmlr-v54-abeille17a), the main idea is to control the regret by showing that randomized exploration ensures a constant probability of optimism. However, compared to the linear bandit analysis, our setting comes with additional challenges: First, due to preference-based learning we require a different regret decomposition accounting for the reference policy. Second, as we observe preference feedback on trajectories rather than policies, we need to apply Freedman’s inequality (see Lemma D.9) to control the deviation between expected and observed features. Lastly, as we assume a PAC RL oracle – in place of an exact maximization oracle – we need to carefully track the resulting approximation error.
@@ 3.2.2 Suboptimality gap
As RPO-Explore collects no preferences during exploration, it may incur linear regret as the policies $\pi_{t}$ can be highly suboptimal. However, Theorem 3.2.2 below shows that the final output policy is $\tilde{\mathcal{O}}(1/\sqrt{T})$-optimal. {restatable}theoremlastiterate For any $\delta\in(0,1)$ and $T\in\mathbb{N}$, Algorithm 2 satisfies, with probability at least $1-\delta$,
@@
Except for the extra $\sqrt{d}$ dependency, which is inherent to approaches based on randomized exploration, [^3] Recently, abeille2025and showed that, in certain cases, the factor $\sqrt{d}$ can be removed by an improved analysis. Their assumptions do not hold in our setting, so whether $\sqrt{d}$ can be avoided remains open. we match the last iterate guarantee proposed by das2024active for a contextual linear bandit setting, but with an algorithm that (i) collects a single batch of preferences and (ii) remains tractable in a full RL setting with trajectory-level feedback, given the PAC RL oracle is tractable.
@@
Algorithm 2 is reminiscent of reward-free RL algorithms (wang2020reward), but in a preference-based setting. To our knowledge, Algorithm 2 is the first tractable algorithm to perform efficient preference-free exploration with trajectory-level preferences, and the first – across preference- or reward-based RL – to do so via randomized exploration. Few other works provide suboptimality guarantees in preference-based RL, but require preference-feedback at every round. wang2023rlhf propose a meta-algorithm interfacing with a PAC-RL oracle that outputs an $\varepsilon$-optimal policy after $\widetilde{\mathcal{O}}(\kappa^{2}d^{3}/\varepsilon^{2})$ queries. A different approach by zhan2024provablereward leverages optimal design to prove a bound of $\widetilde{\mathcal{O}}((|S|^{2}|A|d+\kappa^{2}d^{2})/\varepsilon^{2})$, but their method relies on an intractable maximization oracle. In comparison, Algorithm 2 achieves a bound of $\widetilde{\mathcal{O}}(\kappa d^{3}/\varepsilon^{2})$, improving the dependence on $\kappa$ over both prior results. The extra $\sqrt{d}$ compared to zhan2024provablereward is expected for randomized (rather than optimistic) exploration (pmlr-v54-abeille17a). Finally, note that an online-to-batch conversion yields similar suboptimality bounds for Algorithm 1, but would require preferences at every round (menard2021fast).
@@
The proof of Theorem 3.2.2, presented in Appendix B, builds on das2024active’s suboptimality analysis for the contextual bandits setting. However, to sidestep the intractability of maximizing an exploration bonus over policies, we leverage randomized exploration (pmlr-v54-abeille17a) to ensure a constant probability of optimism. This allows us to derive a bound on the output policy’s suboptimality that mirrors the regret bound, without needing additional assumptions.
@@ 3.3 Practical limitations
While Algorithm 1 is tractable and efficient, it presents certain limitations. Issuing preference queries at each round (line 7) is impractical, due to the need for continuous feedback, and requesting a label for all trajectory pairs can be expensive and inefficient, as many comparisons may be uninformative. Moreover, the large regret of Algorithm 2 may be undesirable for certain applications. The next section introduces a refined regret-minimization algorithm that addresses these issues by decoupling trajectory collection from query selection and querying only the most informative comparisons.
@@ 4.1 Algorithm
Input: $T\in\mathbb{N}$, $\delta\in(0,1)$, $\lambda>1$, $C>0$
1
2Initialize: $\varepsilon=1/\sqrt{T}$; $\delta^{\prime}=\delta/5$; $V_{1}=W_{1}=\lambda I$; $\mathcal{D}_{1}=\mathcal{D}=\emptyset$; $t_{\operatorname{stop}}=1$ ; $\hat{\theta}_{1}=0\,\text{;}\;\pi^{\prime}$.
3for $t=1,2,\dots,T$ do
4 if $\det(W_{t})>(1+C)\det(V_{t_{\operatorname{stop}}})$ then
    $\mathcal{D}_{\text{opt}},V_{t}=\operatorname{D-OptDes}(\mathcal{D},V_{t_{\operatorname{stop}}},W_{t})$
    // Optimal design
5    $\mathcal{D}_{t}=\mathcal{D}_{t_{\operatorname{stop}}}$
@@
       $\mathcal{D}_{t}=\mathcal{D}_{t}\cup\{(x,y)\},x=\phi(\tau)-\phi(\tau^{\prime}),y=\mathbbm{1}(\tau\succ\tau^{\prime})$
       // Preferences
7
@@
    $\hat{\theta}_{t}\in\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t}}(\theta)$
@@
9    $t_{\operatorname{stop}}=t$, $\mathcal{D}=\emptyset$, $\pi^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\hat{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$
@@
 $\tilde{\theta}_{t}\sim\mathcal{N}(\hat{\theta}_{t_{\operatorname{stop}}},\beta_{t_{\operatorname{stop}}}(\delta^{\prime})^{2}V_{t_{\operatorname{stop}}}^{-1})$
@@
 $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$
@@
12 $\mathcal{D}=\mathcal{D}\cup\{(\tau_{t},\tau^{\prime}_{t})\}$, $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$ with $\tau_{t}\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}_{t}\sim\mathbb{P}_{\pi^{\prime}}$
13   If $t=t_{\operatorname{stop}}$, then $W_{t+1}=V_{t}+x_{t}x_{t}^{\top}$, else: $W_{t+1}=W_{t}+x_{t}x_{t}^{\top}$
14
@@
Algorithm 3 LRPO-OD-Regret  (lazy randomized preference optimization with optimal design)
@@
We use an idea from abbasi2011improved to collect many trajectory pairs without querying the preference oracle. The modification compared to Algorithm 1 is collecting trajectories without updating the MLE $\hat{\theta}_{t_{\operatorname{stop}}}$ until the information gain, represented by $\det(V_{t})$, increases by a multiplicative constant; see line 4 of Algorithm 3. We show that this procedure limits the number of batches to $\mathcal{O}(\log(T))$. In other words, the average (over batches) size of a given batch is of order $\mathcal{O}(T/\log(T))$. A key advantage of this lazy update structure is that the preference queries (line 7) can be collected in parallel across all trajectory pairs within a batch. This significantly reduces the workload of the preference oracle, e.g. a human annotator, by eliminating the need for round-by-round feedback, and the algorithm no longer pauses at each timestep to wait for preference labels.
@@
Due to the submodularity of the $\log\det$ function for $\lambda$ greater than one, the greedy procedure of Algorithm 4 achieves an (11/e)(1-1/e)-approximation to the optimal solution; see (nemhauser1978analysis; krause2008near) and Appendix E.1.
Another key feature of Algorithm 4 is its early stopping rule: the while loop terminates early if $\det(V)$ exceeds the determinant of the naive design $W$; if this never happens we simply return $W$. When this early termination is satisfied, Algorithm 3 requires fewer preferences than Algorithm 1. Since the optimal design maximizes the information gain (as measured by $\det(V)$), this termination condition is expected to be satisfied frequently in practice.
Input: Dataset $\mathcal{D}$, previous design matrix $V$, current design matrix $W$ including $\mathcal{D}$.
1 Initialize dataset $\mathcal{D}_{\text{opt}}=\emptyset$
2 while $\det(V)<\det(W)$ and $|\mathcal{D}_{\text{opt}}|\leq|\mathcal{D}|$ do